OpenStudy (anonymous):

what would the followign equation simplify to: x - ( x - 1 - [ x - 2 - ( x - 3)})

OpenStudy (anonymous):

solve the square brackets first. [x-2-x+3] = -1 then the inner brackets x-1-(-1) = x-1+1 = x ; then x-x = 0;

OpenStudy (anonymous):

so, as you solve for each inner brackets, you can remove them?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

x - ( x - 1 - [ x - 2 - ( x - 3)}) x - ( x - 1 - [ x - 2 - x + 3)}) x - ( x - 1 - [ x - x +1)}) x - ( x - 1 - 1) x - ( x - 2) x - x + 2 2

OpenStudy (anonymous):

good luck guys

OpenStudy (anonymous):

Thank you, Mark o. and santoshf